Supplemental Operations Report – January 28th

Date: 1/28/2024
Name of person filing report: Sergii Iakymov
Reason for Report: Routine
Non-Nominal Systems: Intern trailer toilet.

Power system:
Solar: Nominal.
SOC Last 24 hours: Max 100%; Min 63%; Avg 74%.
VDC Last 24 hours: Max 59.47V; Min 43.08; Avg 51.17V.
Generator run time: 4864.5 hours. Serviced on 01222024, oil changed.

Propane Readings:
Station Tank: 80%
Director Tank: 81%
Intern Tank: 84%
Generator Tank: 69%

Hab Static Tank – 331.1 gallons
GreenHab – 131 gallons
Outpost tank – 320 gallons
Science Dome – 0 gallons
Septic Tanks emptied: No

Sojourner rover used: No.
Hours: 196.2
Beginning Charge: 100 %
Ending Charge: 100 %
Currently Charging: Yes
Notes on Rovers: All rovers refilled on 01/25/2024. All rear view mirrors removed from rovers. Three of four mirrors got broken by crews by bumping into them with Eva suits helmets.

ATV: ATV’s Used: None. Nothing to report.

Hab Car used and why, where: To Hanksville for supplies. Need to be serviced.
Crew Car used and why, where: To Hanksville for supplies.
General notes and comments: N/A.

Summary of Internet: All nominal.
EVA suits and radios: All nominal.
Campus wide inspection, if action taken, what and why: All nominal.
Summary of Hab Operations: UV lamp in water filter needs to be replaced, pending. Main heater safety switch replacement pending.
Summary of GreenHab Operations: Heater fixed on 10/26/2024. Control board and thermostat wire replaced
Summary of SciDome Operations: All nominal.
Summary of Observatories Operations: Robotic observatory weather stripping has been resealed with new adhesive 01/26/2024.
Summary of RAM Operations: All nominal.
Summary of Outpost Operations: Intern trailer heater needs to be picked up from Grand Junction. Intern trailer toilet water module assembly is leaking, parts ordered.
Summary of Health and Safety Issues: All nominal.

Copyright © The Mars Society. All rights reserved. | Main Site